# Multimodal Interpretation of Remote Sensing Images: Dynamic Resolution Input Strategy and Multi-scale Vision–Language Alignment Mechanism

### v1

@@ 2.3 Multimodal Understanding of Remote Sensing Images
With the rapid advancement of remote sensing technology, the types and volumes of acquired remote sensing data have expanded explosively, covering optical imagery, Synthetic Aperture Radar (SAR) images, LiDAR point cloud data, hyperspectral images, and datasets from diverse sensors and platforms. These data sources vary in spatial resolution, spectral information, temporal frequency, and imaging principles, but each carries inherent limitations: optical imagery is susceptible to weather and lighting interference; SAR can penetrate clouds and precipitation yet lacks fine spatial resolution and texture details[ref52]; LiDAR provides high-precision 3D terrain information but involves complex processing and higher costs[ref53]. Thus, multimodal remote sensing image understanding, which fuses multi-source, multi-dimensional data, can fully leverage the complementary advantages of each modality, enabling more comprehensive and accurate object recognition and information extraction.
@@
Pixel-level fusion: Aligns and combines data from different modalities at the pixel level to generate composite images containing multimodal information, laying a data foundation for subsequent analysis[ref16]. However, this method requires high spatial alignment accuracy and is typically used for image enhancement and composite feature extraction.
Feature-level fusion: First extracts features independently from each modality, then concatenates or weights these feature vectors[ref54]. In recent years, deep learning approaches (e.g., multi-branch networks, attention mechanisms, cross-modal transformers) have been widely applied here to explore inter-modal relationships and enhance feature representation[ref16].
Decision-level fusion: Performs classification or detection separately on each modality, then integrates results via weighting, voting, or rule-based strategies[ref55]. This approach is simple and flexible, but its performance largely depends on the quality of individual single-modal models.
@@
In recent years, deep learning has driven progress in this field: CNN-based multi-branch architectures extract features from each modality separately, then enable information interaction via fusion layers[ref46]; attention-based methods dynamically adjust modality weights to focus on key regions and critical features[ref44]; Transformer-based multimodal architectures effectively model long-range inter-modal dependencies, boosting feature richness and robustness[ref50]. These methods have promoted multimodal remote sensing applications in land use classification, urban planning, disaster monitoring (e.g., floods, fires), environmental protection, and agricultural assessment, supporting intelligent management across related industries.
In the future, as remote sensing data accumulates and computational technologies advance, multimodal remote sensing fusion is expected to move toward greater intelligence, automation, and real-time processing. By integrating big data analytics, cloud computing, and artificial intelligence, it will elevate remote sensing information services, providing robust technical support for Earth observation, sustainable development, and emergency management[ref56].

### v2

@@ 2.3 Multimodal Understanding of Remote Sensing Images
With the rapid advancement of remote sensing technology, the types and volumes of acquired remote sensing data have expanded explosively, covering optical imagery, Synthetic Aperture Radar (SAR) images, LiDAR point cloud data, hyperspectral images, and datasets from diverse sensors and platforms. These data sources vary in spatial resolution, spectral information, temporal frequency, and imaging principles, but each carries inherent limitations: optical imagery is susceptible to weather and lighting interference; SAR can penetrate clouds and precipitation yet lacks fine spatial resolution and texture details[54]; LiDAR provides high-precision 3D terrain information but involves complex processing and higher costs[42]. Thus, multimodal remote sensing image understanding, which fuses multi-source, multi-dimensional data, can fully leverage the complementary advantages of each modality, enabling more comprehensive and accurate object recognition and information extraction.
@@
Pixel-level fusion: Aligns and combines data from different modalities at the pixel level to generate composite images containing multimodal information, laying a data foundation for subsequent analysis[17]. However, this method requires high spatial alignment accuracy and is typically used for image enhancement and composite feature extraction.
Feature-level fusion: First extracts features independently from each modality, then concatenates or weights these feature vectors[14]. In recent years, deep learning approaches (e.g., multi-branch networks, attention mechanisms, cross-modal transformers) have been widely applied here to explore inter-modal relationships and enhance feature representation[17].
Decision-level fusion: Performs classification or detection separately on each modality, then integrates results via weighting, voting, or rule-based strategies[29]. This approach is simple and flexible, but its performance largely depends on the quality of individual single-modal models.
@@
In recent years, deep learning has driven progress in this field: CNN-based multi-branch architectures extract features from each modality separately, then enable information interaction via fusion layers[5]; attention-based methods dynamically adjust modality weights to focus on key regions and critical features[6]; Transformer-based multimodal architectures effectively model long-range inter-modal dependencies, boosting feature richness and robustness[26]. These methods have promoted multimodal remote sensing applications in land use classification, urban planning, disaster monitoring (e.g., floods, fires), environmental protection, and agricultural assessment, supporting intelligent management across related industries.
In the future, as remote sensing data accumulates and computational technologies advance, multimodal remote sensing fusion is expected to move toward greater intelligence, automation, and real-time processing. By integrating big data analytics, cloud computing, and artificial intelligence, it will elevate remote sensing information services, providing robust technical support for Earth observation, sustainable development, and emergency management[11].